Problém 3-2 (algorithmic) E Question Help Compare the interest earned by $15,000 for five years at 12% simple interest with that earned by the same amount for five years at 12% compounded annually The interest earned by $15,000 for five years at 12% simple interest is S (Round to the nearest dollar )
Answer #1

Interest earned by 15000 for five years at simple interest rate = (15000x12*5)/100= 9000

interest earned by 15000 compounded annually--

Amount after 5 years= 15000(1+0.12)^5 = 26435.125

Compounded interest earned= 26435.125 - 15000= 11435.125